OpenStudy (anonymous):

12. Use the function to answer the question. f(x) = 10x - 5 What is the value of f^–1(–4)? (Points : 4)

OpenStudy (ranga):

f(x) = 10x - 5 For example, if I ask you to find f(6), you will put x = 6 in the equation and come up with: f(6) = 10(6) - 5 = 55. But if I ask you what is f^-1(-4) what it means is what is the value of x that will make the f(x) = -4? f^-1(x) is called the inverse function. Set the left side to -4 and solve for x. f(x) = 10x - 5 -4 = 10x - 5. Solve for x.

OpenStudy (anonymous):

Your "lil bro" should have paid attention in class as these are simple math problems, too elementary; he should also have done his homework before the very last minute. I would have happily helped if you, oh, i mean your "lil bro", had at least attempted to solve the initial portion or simply said I do not understand it. So, let your "lil bro" fail and next time he will pay attention in class and won't put homework until the last moments, literally as it appears. It would be an excellent lesson learned. No pun intended nor any meanness. lol

OpenStudy (anonymous):

I found this on the internet. wut is this

OpenStudy (anonymous):

I know u want 2 laugh jon

OpenStudy (anonymous):

say lol Say lol backpack backpack yay

OpenStudy (anonymous):

CLOSE THIS @JonathanJamesKeeney . Jon is mute Dr. Ranga